# Sparse Actuation for LPV Systems with Full-State Feedback in $\mathcal{H}_{2}/\mathcal{H}_{\infty}$ Framework

### v2

@@ III-B Sparse Actuation for Full State-Feedback LPV Systems
The transfer function matrix from $w$ to $z$ of the closed-loop system in Eq. 2, which is represented by $\mathcal{G}_{z}(\rho,s)$, can be given as
@@
To achieve sparseness in the control input $u$, the $\mathcal{H}_{2}$ norm of the transfer function from $w$ to each $u_{i}$ is minimized. This transfer function, represented by $\mathcal{G}_{u_{i}}(\rho,s)$, corresponds to the $i^{th}$ component of $u=\begin{bmatrix}u_{1}&\cdots&u_{N_{u}}\end{bmatrix}^{\top}$. Using Eq. 1, $\mathcal{G}_{u_{i}}(s)$ can be expressed as,
@@
[…]
@@
$&\min_{\Gamma,K}\|\Gamma\|_{l_{1},\alpha},\\ \text{ subject to }&\left\|\mathcal{G}_{z}(\rho,s)\right\|_{\mathcal{H}_{\infty}}<\gamma_{0},\\ &\left\|\mathcal{G}_{u_{i}}(\rho,s)\right\|_{\mathcal{H}_{2}}<\sqrt{\gamma_{i}},\quad i=1,2,\cdots,N_{u}.\end{aligned}\right\}$ (6)
@@
$&\min_{\Gamma>0,X>0,W}\|\Gamma\|_{l_{1},\alpha}\text{ subject to }\\ &\begin{bmatrix}M_{11}(\rho)&B_{w}(\rho)&\left(C_{z}(\rho)X(\rho)+D_{u}(\rho)W(\rho)\right)^{\top}\\ *&-\gamma_{0}I&D_{w}(\rho)^{\top}\\ *&*&-\gamma_{0}I\end{bmatrix}<0,\\ &M_{11}(\rho)+B_{w}(\rho)B_{w}(\rho)^{\top}<0,\\ &\begin{bmatrix}-\gamma_{i}&\operatorname{row}_{i}(W(\rho))\\ *&-X(\rho)\end{bmatrix}<0,\quad i=1,2,\cdots,N_{u};\end{aligned}\right\}$ (7)
@@
$&\min_{\Gamma>0,X>0,W}\|\Gamma\|_{l_{1},\alpha}\text{ subject to }\\ &\begin{bmatrix}-Z(\rho)&C_{z}(\rho)X(\rho)+D_{u}(\rho)W(\rho)\\ *&-X(\rho)\end{bmatrix}<0,\\ &\operatorname{sym}\left(A(\rho)X(\rho)+B_{u}(\rho)W(\rho)\right)+B_{w}(\rho)B_{w}(\rho)^{\top}<0\text{, }\\ &\operatorname{tr}(Z(\rho))<\gamma_{0}^{2}\\ &\begin{bmatrix}-\gamma_{i}&W(\rho)\\ *&-X(\rho)\end{bmatrix}<0,\quad i=1,2,\cdots,N_{u}\end{aligned}\right\}$ (9)
@@ IV-A Sparse Actuation with $\mathcal{H}_{\infty}$ Full-State Feedback
Here, we address the $\mathcal{H}_{\infty}$ full-state feedback control design aimed at constraining the perturbations in the wing shown in Fig. 1 caused by wind disturbances. Fig. 2 and Fig. 3 present a comparison between the state trajectories of the open-loop and closed-loop simulation. A stable closed-loop performance can be noticed as opposed to the open-loop case. Additionally, significant difference can be seen between the closed-loop response of LTI and LPV models as the LTI models has higher overshoot and settling time.
@@
Fig. 4 and Fig. 5 illustrate the $\|u_{i}(t)\|_{\infty}$ achieved from the $\mathcal{H}_{\infty}$ optimal design for a specific value of $\gamma_{0}$ and varying values of $\gamma_{UB}$. The parameter $\gamma_{UB}$ imposes an upper bound on the control input to emulate hardware limitations. We observe from Fig. 4, that for the controller derived using LPV system mode, actuators 1 is not required to achieve the desired closed-loop disturbance attenuation due to a negligible value of $\|u_{i}(t)\|_{\infty}$. Additionally, the control inputs of actuators 2, 3, and 4 are relatively small. Although the LTI model results in a higher degree of sparseness, the controller performance is subpar as discussed in Fig. 2 and Fig. 3 as the controller does not account for the underlying nonlinearities in the system dynamics..
@@
As more stringent constraints are placed on the upper bound $\gamma_{UB}$, the control distribution adapts accordingly as expected. As observed in Fig. 5, each actuator has an increased contribution of control inputs to meet similar performance requirements as that in the previous case. This highlights the inherent trade-off between actuator sparsity and the maximum allowable control magnitude.
@@ IV-B Sparse Actuation with $\mathcal{H}_{2}$ Full-State Feedback
Here, we address the $\mathcal{H}_{2}$ full-state feedback disturbance rejection problem. Fig. 6 and Fig. 7 present a comparison between the state trajectories of the open-loop and closed-loop simulations. It is observed that the settling time and overshoot for the system is higher than that in the case of $\mathcal{H}_{\infty}$ optimal control. This is because, unlike $\mathcal{H}_{2}$ formulation, $\mathcal{H}_{\infty}$ explicitly considers the worst-case disturbances and provides more robust performance, which inherently results in less oscillatory behavior under external perturbations or uncertainties.
@@
Fig. 8 and Fig. 9 illustrate the $\|u_{i}(t)\|_{\infty}$ achieved from the $\mathcal{H}_{2}$ optimal design for a specific value of $\gamma_{0}$ and varying values of $\gamma_{UB}$. We observe from Fig. 8, for $\mathcal{H}_{2}$ case, that actuators 4, and 5 contribute majorly to the control inputs to achieve the desired closed-loop disturbance attenuation. The $\|u_{i}(t)\|_{\infty}$ for the remaining actuators is significantly lower.
@@
With more stringent constraints imposed on the upper bound $\gamma_{UB}$, the control distribution adjusts accordingly, similar to $\mathcal{H}_{\infty}$ case. As observed in Fig. 9, no additional actuators are required to achieve the same level of performance. However, the contribution of only actuator 4 increases significantly. This is due to the fact that $\mathcal{H}_{2}$ control focuses on minimizing the mean-square response of the system output to disturbances, aiming to reduce the average energy of the output signal. Consequently, while a higher degree of actuation sparsity is observed, it comes at the cost of a longer settling time.
@@ V Conclusions
In conclusion, this paper introduced novel convex optimization formulations for the codesign of actuation architecture and control laws, guaranteeing the specified $\mathcal{H}_{2}/\mathcal{H}_{\infty}$ closed-loop performance in nonlinear systems modeled in LPV form. Detailed in theorem 1 and theorem 2, these formulations were applied to the vibration control of a flexible wing modeled nonlinearly. Our approach could effectively identify a sparse actuation architecture by minimizing the one-norm of actuator magnitude limits while maintaining desired closed-loop performance. This was verified through nonlinear simulations, confirming that actuators with zero magnitude limits could be eliminated without performance loss. Furthermore, our comparisons revealed that sparse solutions based on LTI models were inferior to those employing LPV models, highlighting the importance of incorporating nonlinearities within the design framework.

### v3

@@ III-B Sparse Actuation for Full State-Feedback LPV Systems
The transfer function matrix from $w$ to $z$ of the closed-loop system in Eq.˜2, which is represented by $\mathcal{G}_{z}(\rho,s)$, can be given as
@@
To achieve sparseness in the control input $u$, the $\mathcal{H}_{2}$ norm of the transfer function from $w$ to each $u_{i}$ is minimized. This transfer function, represented by $\mathcal{G}_{u_{i}}(\rho,s)$, corresponds to the $i^{th}$ component of $u=\begin{bmatrix}u_{1}&\cdots&u_{N_{u}}\end{bmatrix}^{\top}$. Using Eq.˜1, $\mathcal{G}_{u_{i}}(s)$ can be expressed as,
@@
$\text{ or }\begin{bmatrix}\sqrt{\gamma_{1}}&\sqrt{\gamma_{2}}&...&\sqrt{\gamma_{N_{u}}}\end{bmatrix}^{\top}=:\sqrt{\Gamma}$
@@
$\left.\begin{aligned} &\min_{\Gamma,K}\|\Gamma\|_{l_{1},\alpha},\\ \text{ subject to }&\left\|\mathcal{G}_{z}(\rho,s)\right\|_{\mathcal{H}_{\infty}}<\gamma_{0},\\ &\left\|\mathcal{G}_{u_{i}}(\rho,s)\right\|_{\mathcal{H}_{2}}<\sqrt{\gamma_{i}},\quad i=1,2,\cdots,N_{u}.\end{aligned}\right\}$ (6)
@@
$\left.\begin{aligned} &\min_{\Gamma>0,X>0,W}\|\Gamma\|_{l_{1},\alpha}\text{ subject to }\\ &\begin{bmatrix}M_{11}(\rho)&B_{w}(\rho)&\left(C_{z}(\rho)X(\rho)+D_{u}(\rho)W(\rho)\right)^{\top}\\ *&-\gamma_{0}I&D_{w}(\rho)^{\top}\\ *&*&-\gamma_{0}I\end{bmatrix}<0,\\ &M_{11}(\rho)+B_{w}(\rho)B_{w}(\rho)^{\top}<0,\\ &\begin{bmatrix}-\gamma_{i}&\operatorname{row}_{i}(W(\rho))\\ *&-X(\rho)\end{bmatrix}<0,\quad i=1,2,\cdots,N_{u};\end{aligned}\right\}$ (7)
@@
$\left.\begin{aligned} &\min_{\Gamma>0,X>0,W}\|\Gamma\|_{l_{1},\alpha}\text{ subject to }\\ &\begin{bmatrix}-Z(\rho)&C_{z}(\rho)X(\rho)+D_{u}(\rho)W(\rho)\\ *&-X(\rho)\end{bmatrix}<0,\\ &\operatorname{sym}\left(A(\rho)X(\rho)+B_{u}(\rho)W(\rho)\right)+B_{w}(\rho)B_{w}(\rho)^{\top}<0\text{, }\\ &\operatorname{tr}(Z(\rho))<\gamma_{0}^{2}\\ &\begin{bmatrix}-\gamma_{i}&W(\rho)\\ *&-X(\rho)\end{bmatrix}<0,\quad i=1,2,\cdots,N_{u}\end{aligned}\right\}$ (9)
@@ IV-A Sparse Actuation with $\mathcal{H}_{\infty}$ Full-State Feedback
Here, we address the $\mathcal{H}_{\infty}$ full-state feedback control design aimed at constraining the perturbations in the wing shown in Fig.˜1 caused by wind disturbances. Fig.˜2 and Fig.˜3 present a comparison between the state trajectories of the open-loop and closed-loop simulation. A stable closed-loop performance can be noticed as opposed to the open-loop case. Additionally, significant difference can be seen between the closed-loop response of LTI and LPV models as the LTI models has higher overshoot and settling time.
@@
Fig.˜4 and Fig.˜5 illustrate the $\|u_{i}(t)\|_{\infty}$ achieved from the $\mathcal{H}_{\infty}$ optimal design for a specific value of $\gamma_{0}$ and varying values of $\gamma_{UB}$. The parameter $\gamma_{UB}$ imposes an upper bound on the control input to emulate hardware limitations. We observe from Fig.˜4, that for the controller derived using LPV system mode, actuators 1 is not required to achieve the desired closed-loop disturbance attenuation due to a negligible value of $\|u_{i}(t)\|_{\infty}$. Additionally, the control inputs of actuators 2, 3, and 4 are relatively small. Although the LTI model results in a higher degree of sparseness, the controller performance is subpar as discussed in Fig.˜2 and Fig.˜3 as the controller does not account for the underlying nonlinearities in the system dynamics..
@@
As more stringent constraints are placed on the upper bound $\gamma_{UB}$, the control distribution adapts accordingly as expected. As observed in Fig.˜5, each actuator has an increased contribution of control inputs to meet similar performance requirements as that in the previous case. This highlights the inherent trade-off between actuator sparsity and the maximum allowable control magnitude.
@@ IV-B Sparse Actuation with $\mathcal{H}_{2}$ Full-State Feedback
Here, we address the $\mathcal{H}_{2}$ full-state feedback disturbance rejection problem. Fig.˜6 and Fig.˜7 present a comparison between the state trajectories of the open-loop and closed-loop simulations. It is observed that the settling time and overshoot for the system is higher than that in the case of $\mathcal{H}_{\infty}$ optimal control. This is because, unlike $\mathcal{H}_{2}$ formulation, $\mathcal{H}_{\infty}$ explicitly considers the worst-case disturbances and provides more robust performance, which inherently results in less oscillatory behavior under external perturbations or uncertainties.
@@
Fig.˜8 and Fig.˜9 illustrate the $\|u_{i}(t)\|_{\infty}$ achieved from the $\mathcal{H}_{2}$ optimal design for a specific value of $\gamma_{0}$ and varying values of $\gamma_{UB}$. We observe from Fig.˜8, for $\mathcal{H}_{2}$ case, that actuators 4, and 5 contribute majorly to the control inputs to achieve the desired closed-loop disturbance attenuation. The $\|u_{i}(t)\|_{\infty}$ for the remaining actuators is significantly lower.
@@
With more stringent constraints imposed on the upper bound $\gamma_{UB}$, the control distribution adjusts accordingly, similar to $\mathcal{H}_{\infty}$ case. As observed in Fig.˜9, no additional actuators are required to achieve the same level of performance. However, the contribution of only actuator 4 increases significantly. This is due to the fact that $\mathcal{H}_{2}$ control focuses on minimizing the mean-square response of the system output to disturbances, aiming to reduce the average energy of the output signal. Consequently, while a higher degree of actuation sparsity is observed, it comes at the cost of a longer settling time.
@@ V Conclusions
In conclusion, this paper introduced novel convex optimization formulations for the codesign of actuation architecture and control laws, guaranteeing the specified $\mathcal{H}_{2}/\mathcal{H}_{\infty}$ closed-loop performance in nonlinear systems modeled in LPV form. Detailed in theorem˜1 and theorem˜2, these formulations were applied to the vibration control of a flexible wing modeled nonlinearly. Our approach could effectively identify a sparse actuation architecture by minimizing the one-norm of actuator magnitude limits while maintaining desired closed-loop performance. This was verified through nonlinear simulations, confirming that actuators with zero magnitude limits could be eliminated without performance loss. Furthermore, our comparisons revealed that sparse solutions based on LTI models were inferior to those employing LPV models, highlighting the importance of incorporating nonlinearities within the design framework.